OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -12, -7, -2, ... @jim_thompson5910

OpenStudy (anonymous):

@ganeshie8

jimthompson5910 (jim_thompson5910):

What's the first term? What's the common difference?

OpenStudy (anonymous):

common difference is 8 @jim_thompson5910

OpenStudy (anonymous):

an = 8 + -2(n - 1)? or is it n+1?

jimthompson5910 (jim_thompson5910):

First term: a1 = -17 Common Difference: d = 5 ------------------------------------------------------- an = a1 + d(n-1) an = -17 + 5(n-1) an = ???